OpenStudy (anonymous):

Two points A and B lie on opposite sides of a river. Another point C is located on the same side of the river as B at a distance of 240 ft from B. If angle ABC is 120° and angle ACB is 15°, find the distance across the river from A to B. (Round your answer to two decimal places.)

jimthompson5910 (jim_thompson5910):

Hint:  use the law of sines to solve for x

jimthompson5910 (jim_thompson5910):

you'll need angle A first, so A+B+C = 180 A+120+15=180 A+135=180 A = 180-135 A = 45